What is the equation of the line that passes through (-2,0) and is perpendicular to y=3x+1?
Answer by mananth(16949)

1 y = 3 x + 1
Divide by 1
y = 3 x + 1
Compare this equation with y=mx+b
slope m = 3.00

The slope of a line perpendicular to the above line will be the negative reciprocal -0.33
m1*m2=-1
The slope of the required line will be -0.33 =-1/3

m= - 1/3 ,point (-2,0)
Find b by plugging the values of m & the point in
y=mx+b
0 = 2/3 + b
b= - 2/3
m= - 1/3
The required equation is y=(-1/ 3)x- 2/ 3
m.ananth@hotmail.ca